# Performance Trade-off and Joint Waveform Design for MIMO-OFDM DFRC Systems

### v1

@@ VI Conclusion
Considering the communication and radar detection performance trade-off, this paper proposed two waveform design strategies for the DFRC system based on MIMO and OFDM technologies in the frequency-selective multipath fading environment. ISI and achievable rate were employed as the performance metrics in the two proposed design strategies, respectively. In the CRB and ISI minimization based waveform design strategy, which can be viewed as an extension of the traditional single carrier based DFRC waveform to the frequency-selective multipath fading environment, we first solved the radar-strict problem and then allowed a tolerable deviation between the actual precoding design and the radar-strict design to achieve the performance trade-off between the radar system and the communication system. In the CRB minimization and achievable rate maximization based waveform design strategy, we proposed the WF-SQP algorithm to solve the optimization problem. Simulation results showed the superiority of the proposed waveform design strategies.

### v2

@@ VI Conclusion
Considering the communication and radar detection performance trade-off, this paper proposed two waveform design schemes for the DFRC system based on MIMO and OFDM technologies in the frequency-selective multipath fading environment. ISI and achievable rate were employed as the performance metrics in the two proposed design schemes, respectively. In the CRB and ISI minimization based waveform design scheme, which can be viewed as an extension of the traditional single carrier based DFRC waveform to the frequency-selective multipath fading environment, we first solved the radar-strict problem and then allowed a tolerable deviation between the actual precoding design and the radar-strict design to achieve the performance trade-off between the radar system and the communication system. In the CRB minimization and achievable rate maximization based waveform design scheme, we proposed the WF-SQP algorithm to solve the optimization problem. Simulation results showed the superiority of the proposed waveform design schemes.